Fermi National Laboratory


Accelerator Update for June 15, 2005 - June 17, 2005

Wednesday June 15
The day shift began with the Tevatron (TeV) cooldown continuing, and with MiniBooNE and NuMI taking beam.

The Antiproton Source (Pbar) resumed stacking around 8 AM.

A Switchyard (SY) expert began tuning beam to the Meson MCenter experiment at 5:30 PM.

At 10:34 PM, the I- Source chopper and a Linac RF station (KRF1) tripped off. The chopper trip was due to a vacuum ion gauge trip and KRF1 tripped off due to a LCW flow switch problem. Both devices reset okay.

Thursday June 16
The midnight shift began with the TeV cooldown continuing, with Pbar stacking, and with MiniBooNE and NuMI taking beam.

At 12:48 AM, a Meson quadrupole power supply (MC2Q2) tripped off and wouldn't reset. An expert adjusted the tune to make up for the missing quadrupole.

At 5:44 AM, Operations reported that the TeV turn on was going slowly due to problems with a vacuum valve in between TeV sectors A2 and A3 and with a dipole power supply (HF11). Operators changed out the regulator for HF11. Vacuum technicians managed to manually move the valve.

Operations put the Recycler into shot setup at 6:39 AM to stash antiprotons.

Operations put the TeV through a dry squeeze at 9:09 AM.

A Meson expert began tuning beam to MCenter and MTest at 1:21 PM.

At 1:29 PM, cryo system technicians reported that the flywheel for a TeV sector E2 wet engine was wobbling and needed to be replaced. They began the work.

At 1:41 PM, the experimenters at D-Zero (D0) reported that they had a power supply failure in their collision hall and needed an access to repair it. Operators prepared the area for access.

At 3:36 PM, a cryo expert reported that the TeV sector E cryo system looked bad.

The D0 access ended and Operations put the TeV through another dry squeeze at 3:50 PM.

A TeV expert began conducting studies at 4:32 PM.

Two Pbar experts began a study period 4:35 PM.

Operations reported a Pbar emittance blowup at 4:50 PM. Stacking was halted while Pbar experts investigated. Stacking resumed at 5:21 PM.

Operations made D0 ready for an access at 8:06 PM. They came out of access about 1 hour and 45 minutes later.

Operations put the TeV through one more dry squeeze at 10:11 PM before going into a shot setup.

Operations put the TeV into shot setup at 10:46 PM.

Friday June 17
The midnight shift began with the TeV in shot setup, and with MiniBooNE taking beam.

Operations reported at 12:55 AM that all protons and antiprotons had been loaded and that they were moving collimators and initiating collisions.

At 1:06 AM, the TeV suffered a nine-house quench during scraping. (Much later, experts reported that a sector C4 vertical separator had sparked during the scrape and this was the root cause of the quenches.)

Operations established beam to MCenter and MTest at 2:29 AM.

Operations put the TeV Separators in conditioning mode at 3:48 AM.

Operations prepared the Collider Detector at Fermilab (CDF) for access at 5:43 AM. CDF came out of access at 7:24 AM.

Operations put the TeV through a dry squeeze at 7:53 AM.

The Plan for Friday and the Weekend
The plan for today is to establish a store this morning. The plans for the weekend are to stack, stash, and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/17/2005   email Fermilab